# OpenPBR: Novel Features and Implementation Details

### v1

@@ 3.1.2 Subsurface
While designing OpenPBR, we considered and experimented with a number of different options. One of the options we tentatively finalized involved setting the average distance that light visibly travels along the surface for each color channel after all scattering and absorption has occurred. We based this on the theory presented in [Christensen2015]. Effectively the user would be controlling the resultant diffusion profile directly (even when simulated using path-traced volumetric scattering). This was conceptually appealing, especially since it aligned well with the perceptual multiple-scattering definition of the subsurface color parameter described below.
@@
A well-known approximation is due to van de Hulst [Kulla2017, HHGTMS], which assumes an index-matched boundary (i.e., $\mathbf{E}_{\mathrm{spec}}=0$, thus $\mathbf{E}_{\mathrm{multi-scatter}}=\mathbf{C}$). According to this, the total reflectance (i.e., multi-scatter albedo) $\mathbf{C}$ from a semi-infinite isotropic medium with no Fresnel (vacuum boundary) and diffusion illumination is given by
@@
This provides the desired single-scatter albedo given multi-scatter albedo $\mathbf{C}$ as $\boldsymbol{\alpha}^{\star}(\mathbf{C})=1-\mathbf{s}^{2}(\mathbf{C})$. [ChiangKutz2016] found their own fitting function for multi-scatter albedo $\boldsymbol{\alpha}^{\star}(\mathbf{C})$, which was later extended by [Hyperion], taking into account TIR from an interface assumed to have IOR 1.4, producing
@@
These formulas are necessarily approximate, as for example they neglect certain effects such as the roughness of the dielectric boundary. Various other alternative analytical formulas which attempt to account more accurately for all the relevant effects are presented by [HHGTMS].
To account for anisotropy, the standard approach is to use the following “similarity relation” [Zhao2014, Hyperion] between the scattering coefficients in media with different anisotropies $g$ and $g^{\star}$:
@@
Note that the similarity relation described above can also be used to make path-traced volume rendering more efficient by reducing the number of scattering events. This can be done for a volume with anisotropy $g$ by setting $g^{\star}$ to any value between $g$ and 0 (0 for isotropic scattering), computing $\boldsymbol{\mu}^{\star}_{s}$ using Equation 49, and simply setting $\boldsymbol{\mu}^{\star}_{a}=\boldsymbol{\mu}_{a}$ (since the distance traveled along the path and therefore the amount of absorption is essentially unchanged). As proposed by [Hyperion], it is even possible to gradually transition from full anisotropic scattering to reduced isotropic scattering over the first several bounces along a path by setting $g^{\star}$ differently at each bounce, gradually reducing it from $g$ to 0. This can allow an implementation to achieve the best of both worlds of visuals and performance: view-dependent, low-order scattering and efficient high-order scattering.
@@
Note that the default value of subsurface_radius_scale is set at $(1,0.5,0.25)$ to approximate the effect of Rayleigh scattering (or Tyndall scattering). If we roughly approximate the wavelength corresponding to each of the RGB channels as $\lambda/\mathrm{nm}\approx 650,550,450$ and assume the radii scale like the reciprocal of the extinction, then since in Rayleigh scattering the extinction scales like $\lambda^{-4}$ for light of wavelength $\lambda$, the expected relative magnitudes of the radii are $(1,(550/650)^{4},(450/650)^{4})\approx(1,0.5,0.25)$, hence the chosen default. This provides a slightly more realistic default look for the subsurface than resulting from gray radii (Figure 22). We considered a more explicit control for providing a Rayleigh-scattering effect (such as the model from [Kutz2021]) but felt that this could always in principle be built of top of the model, and having the default look be reminiscent of Rayleigh scattering was sufficient.
@@ 3.1.3 Translucent base
The default 20 was chosen since common real materials with the highest dispersion have Abbe numbers roughly greater than or equal to this (for example, different types of glass have Abbe numbers between 20 and 91, water and diamond both have Abbe numbers between 55 and 56, while the oxide mineral rutile, composed of titanium dioxide, exhibits extremely high dispersion with an Abbe number of 9.87 [Polyanskiy2023]).
@@
Note that dispersion can cause banding artifacts and color noise if not properly managed, however these issues can be mitigated using a careful strategy of wavelength sampling, color-channel selection, and MIS. For sampling wavelengths and incorporating spectral effects inside an RGB renderer, multiple approaches exist, such as the empirical one described by [Kutz2021]. The wavelengths from such a “locally spectral” approach can also be used for rendering thin-film iridescence, as mentioned in Section 4. For selecting a color channel to use for the refracted ray direction, it is useful to multiply the path throughput into the probabilities. This way, the channel carrying the most energy along the path so far will be preferentially used for continuing the path. This keeps the magnitude of the path throughput balanced and avoids generating paths that don’t contribute to the image.
For example, if one scattering event selects the red channel and generates a refracted path that is only valid for the red channel, the next scattering event will also use the red channel. While careful color-channel selection can reduce variance, it doesn’t directly target color noise. Color noise is inevitable when individual paths are only valid for a single wavelength, however when dispersion occurs on a rough microfacet surface, the resultant transmission lobes can overlap significantly, so paths can be shared across color channels using MIS (as described by [Kutz2021]). Similar approaches can also be leveraged to reduce noise when using path tracing to render volumes with wavelength-varying extinction coefficients: specifically, the path throughput (along with the single-scattering albedo) can be incorporated into the color-channel selection probabilities for distance sampling, and paths can be shared among color channels using MIS (as described by [ChiangKutz2016] and [Kutz2017]).
@@ 3.2 Metallic base: the F82-tint model
Metals are completely opaque and have a characteristic and familiar form of specularity due to the Fresnel factor for conductors differing from that of dielectrics. The conductor Fresnel reflection curve is parameterized by the colors at normal and near-grazing incidence (base_color and specular_color respectively, scaled by the corresponding weights). This allows for art-directable variation in reflectivity toward the grazing edges by directly specifying the (texturable) colors at normal and grazing incidence to simulate the dip in reflectivity observed in real metals, or just for artistic effect [Hoffman2019]. Note that these two color parameters are also used for the non-metallic (i.e., dielectric) specular and diffuse BRDFs.
@@
The specific model we stipulate for the metallic Fresnel factor $\mathbf{F}_{\mathrm{metal}}(\mu)$ is the “F82-tint” model of [Kutz2021], which extends previous work by [Hoffman2019]. This is based on the standard Schlick approximation to the metallic Fresnel factor, where $\mathbf{F}_{0}$ is the RGB reflectivity at normal incidence (i.e., $\mathtt{base\_weight}*\mathtt{base\_color}$), and $\mu$ is the cosine of the incident angle:
@@
Note that the clamp at the lower end is applied since $\mathbf{F}_{82}(\mu)$ can become negative for some values of $\mu$, which is not physically meaningful [Hoffman2019]. Note that the edge cannot be brighter than the standard Schlick term, but this is generally true in real metals. We consider this a benefit of this parameterization, since it makes it impossible to produce physically implausible metals with excessively bright edges.
Figure 25 shows renders of chromium, lead, zinc, gold and copper, with and without the physically correct specular_color F82-tint color. In the cases of chromium and lead, the effect of the F82-tint color is visually significant near the edge. In the case of zinc it is less obvious, and barely perceptible in the case of gold and copper. In Appendix D, we provide a table of computed fits to the F82-tint model for a variety of real metals, using publicly available tabulated IOR data [Polyanskiy2023].
@@
The use of base_color and specular_color to control the reflectivity of both the metallic and dielectric base is a convenience that originates from the model of [Burley2012], and corresponds to the popular “metalness” workflow where a single set of textures, in conjunction with a metalness map, can be used to define both metallic and non-metallic regions on the same object.
@@ 3.4.1 Hazy specular
A strongly requested feature is a more configurable microfacet NDF which can support a “hazy gloss” look. The coat layer does not provide equivalent functionality, since the coat produces a strong Fresnel effect near grazing, roughens the underlying base, and generates darkening due to inter-reflections. A number of models for a generalized microfacet NDF have been proposed. [Ribardiere2017] introduced a BRDF based on the Student’s t-distribution. [Burley2012] introduced the “generalized Trowbridge-Reitz” (GTR) NDF. These add a 1-parameter control over the NDF shape. [Barla2018] introduced a model with a blend of a primary lobe and secondary lobe differing in roughness, expressed as a sum of a “specular core” and a “surrounding halo”, in such a way that the halo width can be adjusted independently of the core brightness.
An alternative, simpler approach based on [Barla2018] that we favour is to simply have the BSDF be a blend between microfacet BSDFs with independent roughnesses, the original “core” lobe and an added “haze” lobe with some mix weight $w_{h}\in[0,1]$ and effective roughness $r_{h}$ greater than the core roughness. We noted that this model is actually more expressive and physically plausible than the 1-parameter models, since the mix weight $w_{h}$ effectively controls the ratio of the energy in the haze relative to the core, while the haze roughness $r_{h}$ can be varied completely independently of this (supporting weak and strong haze, with independent low or high added roughness).
@@ 3.4.2 Retro-reflection
Rendering of retro-reflective materials is useful in a variety of contexts (typically for safety applications such as road markings, signs, vehicles and clothing). Such materials are usually designed to be retro-reflective via a substructure of elements that preferentially scatter light backwards. A number of CG models have been proposed ([Belcour2014, Guo2017, Guo2018b]). For practical purposes in visual effects, we are interested in a model which is visually plausible at least, provably energy-conserving, efficient and easy to implement, and ideally a small modification to well-understood microfacet models so it can be implemented as a modification to the specular lobe.
A model we are evaluating is based on the previously published “back-vector” formulation [Belcour2014] which meets these requirements. This empirically based approach simply replaces the half-vector $\mathbf{h}$ in the microfacet model with the back-vector $\mathbf{b}$ defined as
@@
[Belcour2014] demonstrated that this formulation achieves a reasonable match to measured retro-reflective materials. Given an implementation of a regular microfacet BRDF, extending it to retro-reflection is then extremely straightforward:
  • Evaluation merely needs to replace $\mathbf{v}$ with $\mathbf{v^{\prime}}$ upfront. Similarly, importance sampling of $\mathbf{l}$ given $\mathbf{v}$ can be realized by replacing $\mathbf{v}$ with $\mathbf{v^{\prime}}$ upfront and then importance sampling the regular microfacet BRDF. This may include low variance sampling using the domain of visible microfacets [Heitz2018].
  • As the albedos of the standard BRDF and retro-reflective BRDF are essentially identical, compensating for energy loss in the sense of [Kelemen2001] can be realized using the same data tables.
@@
The currently recommended thin-film model is that of [Belcour2017], which pre-integrates interference effects using Fourier-domain convolutions and Gaussian filtering. This method efficiently produces high-quality fringe patterns in an RGB rendering context, but it can be challenging to implement and may introduce inaccuracies in some cases as it assumes that Fresnel amplitude and phase coefficients remain constant across each spectral band, which limits the model’s ability to capture wavelength-dependent dispersion effects.
@@
The same wavelengths can also be reused to model dispersion, while all other BSDF components are free to ignore them and operate in RGB as usual. This approach uses only the Airy summation from Belcour and Barla (Equation 3 from [Belcour2017]) – a method whose origins trace back to 19th-century studies of thin-plate interference by G.B. Airy and others – but requires additional per-wavelength computations and assembling the necessary formulas from multiple sources rather than a single reference. We attempt to collect the necessary equations and implementation instructions in one place in Appendix E.
@@
  • In the case of the dielectric base, the thin film should also generate color fringes in the transmission lobe. This is important for example when rendering soap bubbles (see [Belcour2017]).
  • In the case of the metallic base, the physics is somewhat ambiguous since the Fresnel factor for metal is defined according to the Schlick-based “F82-tint” parameterization (as described in Section 3.2), which does not specify the underlying physical complex IOR. We suggest here that some reasonable approximation is employed to map the Fresnel factor to the best matching effective complex IOR, for example that described by [Gulbrandsen2014].
@@ 5.1 Darkening
To describe what we mean by “counteract the darkening”, we write the (physically darkened) coat albedo at normal incidence in the general form
$\mathbf{E}_{c}=F_{0}+(1-F_{0})\mathbf{E}^{\prime}_{c}\ ,$ (73)
where $F_{0}$ is the normal incidence Fresnel factor of the coat, and $\mathbf{E}^{\prime}_{c}$ represents the albedo due to transmission into the coat medium (and scattering off the base substrate, potentially multiple internal reflections off the coat interface, and re-transmission back out). We then require that the effect of $\delta=\mathtt{coat\_darkening}$ is to multiply $\mathbf{E}^{\prime}_{c}$ by an RGB boost factor
$\mathbf{B}(\delta)=\mathrm{lerp}\biggl(\frac{\mathbf{T}_{\mathrm{coat}}^{2}\mathbf{E}_{b}}{\mathbf{E}^{\prime}_{c}},1,\delta\biggr)\ ,$ (74)
@@
$\mathbf{E}_{c}=F_{0}+(1-F_{0})\mathbf{E}^{\prime}_{c}\mathbf{B}(\delta)\;\rightarrow\;F_{0}+\mathbf{T}^{2}_{\mathrm{coat}}\,\mathbf{E}_{b}(1-F_{0})$ (75)
@@
Note that we have not specified the detailed physical mechanism by which the boosting of $\mathbf{E}^{\prime}_{c}$ occurs. Instead, we have only defined what the resultant effective albedo modification (at normal incidence) has to be as coat_darkening is varied. While this defines the behavior at the level of the required albedos, in practice implementations will need to develop some specific approximation for the coat darkening effect consistent with their physical approximations. Due to the physical effect of the darkening due to the internal reflections, the base BSDF would generally need to be altered in order for the coated base to achieve the un-darkened albedo. While this can be implementation dependent, we suggest here a reasonably simple, efficient scheme that captures the essential behavior.
@@
float C = P.coat_weight;
@@
vec3 Cb = P.base_weight * P.base_color;
@@
float Sc = P.subsurface_color;
vec3 WHITE = vec3(1.f);
@@
vec3 E_dielec = mix(mix(S, Cb, Sc), vec3(1.f-F0), T); // dielectric base albedo
vec3 E_metal = clamp(Cb * xi, 0.f, 1.f); // metallic base albedo
@@
vec3 Delta = max(1.f - K, 0.f) / (WHITE - E_base*K*Cc); // base darkening factor, Eqn.76
@@
// Apply albedo scaling to model the layering of the coat on the base,
// taking into account the base darkening. First, the coat BSDF has the
// coat_weight presence weight functioning as a multiplier:
@@
// The base BSDFs are then multiplied by the coat_throughput factor:
@@
vec3 coat_throughput = mix(WHITE, base_darkening * (WHITE - coat_albedo), C)

### v2

@@ 3.1.2 Subsurface
While designing OpenPBR, we considered and experimented with a number of different options. One of the options we tentatively finalized involved setting the average distance that light visibly travels along the surface for each color channel after all scattering and absorption has occurred. We based this on the theory presented in [12]. Effectively the user would be controlling the resultant diffusion profile directly (even when simulated using path-traced volumetric scattering). This was conceptually appealing, especially since it aligned well with the perceptual multiple-scattering definition of the subsurface color parameter described below.
@@
A well-known approximation is due to van de Hulst [35, 13], which assumes an index-matched boundary (i.e., $\mathbf{E}_{\mathrm{spec}}=0$, thus $\mathbf{E}_{\mathrm{multi-scatter}}=\mathbf{C}$). According to this, the total reflectance (i.e., multi-scatter albedo) $\mathbf{C}$ from a semi-infinite isotropic medium with no Fresnel (vacuum boundary) and diffusion illumination is given by
@@
This provides the desired single-scatter albedo given multi-scatter albedo $\mathbf{C}$ as $\boldsymbol{\alpha}^{\star}(\mathbf{C})=1-\mathbf{s}^{2}(\mathbf{C})$. [11] found their own fitting function for multi-scatter albedo $\boldsymbol{\alpha}^{\star}(\mathbf{C})$, which was later extended by [9], taking into account TIR from an interface assumed to have IOR 1.4, producing
@@
These formulas are necessarily approximate, as for example they neglect certain effects such as the roughness of the dielectric boundary. Various other alternative analytical formulas which attempt to account more accurately for all the relevant effects are presented by [13].
To account for anisotropy, the standard approach is to use the following “similarity relation” [58, 9] between the scattering coefficients in media with different anisotropies $g$ and $g^{\star}$:
@@
Note that the similarity relation described above can also be used to make path-traced volume rendering more efficient by reducing the number of scattering events. This can be done for a volume with anisotropy $g$ by setting $g^{\star}$ to any value between $g$ and 0 (0 for isotropic scattering), computing $\boldsymbol{\mu}^{\star}_{s}$ using Equation 49, and simply setting $\boldsymbol{\mu}^{\star}_{a}=\boldsymbol{\mu}_{a}$ (since the distance traveled along the path and therefore the amount of absorption is essentially unchanged). As proposed by [9], it is even possible to gradually transition from full anisotropic scattering to reduced isotropic scattering over the first several bounces along a path by setting $g^{\star}$ differently at each bounce, gradually reducing it from $g$ to 0. This can allow an implementation to achieve the best of both worlds of visuals and performance: view-dependent, low-order scattering and efficient high-order scattering.
@@
Note that the default value of subsurface_radius_scale is set at $(1,0.5,0.25)$ to approximate the effect of Rayleigh scattering (or Tyndall scattering). If we roughly approximate the wavelength corresponding to each of the RGB channels as $\lambda/\mathrm{nm}\approx 650,550,450$ and assume the radii scale like the reciprocal of the extinction, then since in Rayleigh scattering the extinction scales like $\lambda^{-4}$ for light of wavelength $\lambda$, the expected relative magnitudes of the radii are $(1,(550/650)^{4},(450/650)^{4})\approx(1,0.5,0.25)$, hence the chosen default. This provides a slightly more realistic default look for the subsurface than resulting from gray radii (Figure 22). We considered a more explicit control for providing a Rayleigh-scattering effect (such as the model from [37]) but felt that this could always in principle be built of top of the model, and having the default look be reminiscent of Rayleigh scattering was sufficient.
@@ 3.1.3 Translucent base
The default 20 was chosen since common real materials with the highest dispersion have Abbe numbers roughly greater than or equal to this (for example, different types of glass have Abbe numbers between 20 and 91, water and diamond both have Abbe numbers between 55 and 56, while the oxide mineral rutile, composed of titanium dioxide, exhibits extremely high dispersion with an Abbe number of 9.87 [42]).
@@
Note that dispersion can cause banding artifacts and color noise if not properly managed, however these issues can be mitigated using a careful strategy of wavelength sampling, color-channel selection, and MIS. For sampling wavelengths and incorporating spectral effects inside an RGB renderer, multiple approaches exist, such as the empirical one described by [37]. The wavelengths from such a “locally spectral” approach can also be used for rendering thin-film iridescence, as mentioned in Section 4. For selecting a color channel to use for the refracted ray direction, it is useful to multiply the path throughput into the probabilities. This way, the channel carrying the most energy along the path so far will be preferentially used for continuing the path. This keeps the magnitude of the path throughput balanced and avoids generating paths that don’t contribute to the image.
For example, if one scattering event selects the red channel and generates a refracted path that is only valid for the red channel, the next scattering event will also use the red channel. While careful color-channel selection can reduce variance, it doesn’t directly target color noise. Color noise is inevitable when individual paths are only valid for a single wavelength, however when dispersion occurs on a rough microfacet surface, the resultant transmission lobes can overlap significantly, so paths can be shared across color channels using MIS (as described by [37]). Similar approaches can also be leveraged to reduce noise when using path tracing to render volumes with wavelength-varying extinction coefficients: specifically, the path throughput (along with the single-scattering albedo) can be incorporated into the color-channel selection probabilities for distance sampling, and paths can be shared among color channels using MIS (as described by [11] and [36]).
@@ 3.2 Metallic base: the F82-tint model
Metals are completely opaque and have a characteristic and familiar form of specularity due to the Fresnel factor for conductors differing from that of dielectrics. The conductor Fresnel reflection curve is parameterized by the colors at normal and near-grazing incidence (base_color and specular_color respectively, scaled by the corresponding weights). This allows for art-directable variation in reflectivity toward the grazing edges by directly specifying the (texturable) colors at normal and grazing incidence to simulate the dip in reflectivity observed in real metals, or just for artistic effect [30]. Note that these two color parameters are also used for the non-metallic (i.e., dielectric) specular and diffuse BRDFs.
@@
The specific model we stipulate for the metallic Fresnel factor $\mathbf{F}_{\mathrm{metal}}(\mu)$ is the “F82-tint” model of [37], which extends previous work by [30]. This is based on the standard Schlick approximation to the metallic Fresnel factor, where $\mathbf{F}_{0}$ is the RGB reflectivity at normal incidence (i.e., $\mathtt{base\_weight}*\mathtt{base\_color}$), and $\mu$ is the cosine of the incident angle:
@@
Note that the clamp at the lower end is applied since $\mathbf{F}_{82}(\mu)$ can become negative for some values of $\mu$, which is not physically meaningful [30]. Note that the edge cannot be brighter than the standard Schlick term, but this is generally true in real metals. We consider this a benefit of this parameterization, since it makes it impossible to produce physically implausible metals with excessively bright edges.
Figure 25 shows renders of chromium, lead, zinc, gold and copper, with and without the physically correct specular_color F82-tint color. In the cases of chromium and lead, the effect of the F82-tint color is visually significant near the edge. In the case of zinc it is less obvious, and barely perceptible in the case of gold and copper. In Appendix D, we provide a table of computed fits to the F82-tint model for a variety of real metals, using publicly available tabulated IOR data [42].
@@
The use of base_color and specular_color to control the reflectivity of both the metallic and dielectric base is a convenience that originates from the model of [10], and corresponds to the popular “metalness” workflow where a single set of textures, in conjunction with a metalness map, can be used to define both metallic and non-metallic regions on the same object.
@@ 3.4.1 Hazy specular
A strongly requested feature is a more configurable microfacet NDF which can support a “hazy gloss” look. The coat layer does not provide equivalent functionality, since the coat produces a strong Fresnel effect near grazing, roughens the underlying base, and generates darkening due to inter-reflections. A number of models for a generalized microfacet NDF have been proposed. [44] introduced a BRDF based on the Student’s t-distribution. [10] introduced the “generalized Trowbridge-Reitz” (GTR) NDF. These add a 1-parameter control over the NDF shape. [4] introduced a model with a blend of a primary lobe and secondary lobe differing in roughness, expressed as a sum of a “specular core” and a “surrounding halo”, in such a way that the halo width can be adjusted independently of the core brightness.
An alternative, simpler approach based on [4] that we favour is to simply have the BSDF be a blend between microfacet BSDFs with independent roughnesses, the original “core” lobe and an added “haze” lobe with some mix weight $w_{h}\in[0,1]$ and effective roughness $r_{h}$ greater than the core roughness. We noted that this model is actually more expressive and physically plausible than the 1-parameter models, since the mix weight $w_{h}$ effectively controls the ratio of the energy in the haze relative to the core, while the haze roughness $r_{h}$ can be varied completely independently of this (supporting weak and strong haze, with independent low or high added roughness).
@@ 3.4.2 Retro-reflection
Rendering of retro-reflective materials is useful in a variety of contexts (typically for safety applications such as road markings, signs, vehicles and clothing). Such materials are usually designed to be retro-reflective via a substructure of elements that preferentially scatter light backwards. A number of CG models have been proposed ([6, 21, 20]). For practical purposes in visual effects, we are interested in a model which is visually plausible at least, provably energy-conserving, efficient and easy to implement, and ideally a small modification to well-understood microfacet models so it can be implemented as a modification to the specular lobe.
A model we are evaluating is based on the previously published “back-vector” formulation [6] which meets these requirements. This empirically based approach simply replaces the half-vector $\mathbf{h}$ in the microfacet model with the back-vector $\mathbf{b}$ defined as
@@
[6] demonstrated that this formulation achieves a reasonable match to measured retro-reflective materials. Given an implementation of a regular microfacet BRDF, extending it to retro-reflection is then extremely straightforward:
  • Evaluation merely needs to replace $\mathbf{v}$ with $\mathbf{v^{\prime}}$ upfront. Similarly, importance sampling of $\mathbf{l}$ given $\mathbf{v}$ can be realized by replacing $\mathbf{v}$ with $\mathbf{v^{\prime}}$ upfront and then importance sampling the regular microfacet BRDF. This may include low variance sampling using the domain of visible microfacets [28].
  • As the albedos of the standard BRDF and retro-reflective BRDF are essentially identical, compensating for energy loss in the sense of [33] can be realized using the same data tables.
@@
The currently recommended thin-film model is that of [5], which pre-integrates interference effects using Fourier-domain convolutions and Gaussian filtering. This method efficiently produces high-quality fringe patterns in an RGB rendering context, but it can be challenging to implement and may introduce inaccuracies in some cases as it assumes that Fresnel amplitude and phase coefficients remain constant across each spectral band, which limits the model’s ability to capture wavelength-dependent dispersion effects.
@@
The same wavelengths can also be reused to model dispersion, while all other BSDF components are free to ignore them and operate in RGB as usual. This approach uses only the Airy summation from Belcour and Barla (Equation 3 from [5]) – a method whose origins trace back to 19th-century studies of thin-plate interference by G.B. Airy and others – but requires additional per-wavelength computations and assembling the necessary formulas from multiple sources rather than a single reference. We attempt to collect the necessary equations and implementation instructions in one place in Appendix E.
@@
  • In the case of the dielectric base, the thin film should also generate color fringes in the transmission lobe. This is important for example when rendering soap bubbles (see [5]).
  • In the case of the metallic base, the physics is somewhat ambiguous since the Fresnel factor for metal is defined according to the Schlick-based “F82-tint” parameterization (as described in Section 3.2), which does not specify the underlying physical complex IOR. We suggest here that some reasonable approximation is employed to map the Fresnel factor to the best matching effective complex IOR, for example that described by [19].
@@ 5.1 Darkening
To define what we mean by “counteract the darkening”, we write the (physically darkened) coat albedo at normal incidence in the general form, similar to Equation 39 and Equation 43:
$\mathbf{E}_{c}=\mathbf{E}_{\mathrm{spec}}+(1-\mathbf{E}_{\mathrm{spec}})\mathbf{E}^{\prime}_{c}$ (73)
where $\mathbf{E}_{\mathrm{spec}}$ is the normal-direction reflectance of all energy reflected from the coat’s dielectric interface without macroscopic transmission, and $\mathbf{E}^{\prime}_{c}$ represents the albedo due to transmission into the coat medium (and scattering off the base substrate, potentially multiple internal reflections off the coat interface, and re-transmission back out). We then require that the effect of $\delta=\mathtt{coat\_darkening}$ is to multiply $\mathbf{E}^{\prime}_{c}$ by an RGB boost factor
$\mathbf{B}(\delta)=\mathrm{lerp}\biggl(\frac{\mathbf{T}_{\mathrm{coat}}^{2}\mathbf{E}_{b}}{\mathbf{E}^{\prime}_{c}},1,\delta\biggr)$ (74)
@@
$\mathbf{E}_{c}=\mathbf{E}_{\mathrm{spec}}+(1-\mathbf{E}_{\mathrm{spec}})\mathbf{E}^{\prime}_{c}\mathbf{B}(\delta)\;\rightarrow\;\mathbf{E}_{\mathrm{spec}}+\mathbf{T}^{2}_{\mathrm{coat}}\,\mathbf{E}_{b}(1-\mathbf{E}_{\mathrm{spec}})$ (75)
@@
Note that we have not specified the detailed physical mechanism by which the boosting of $\mathbf{E}^{\prime}_{c}$ occurs. Instead, we have only defined what the resultant effective albedo modification (at normal incidence) has to be as coat_darkening is varied. While this defines the behavior at the level of the required albedos, in practice implementations will need to develop some specific approximation for the coat darkening effect consistent with their physical approximations. Due to the physical effect of the darkening due to the internal reflections, the base BSDF would generally need to be altered (most simply by multiplying by a boost factor) in order for the coated base to achieve the un-darkened albedo. While this can be implementation dependent, we suggest here a reasonably simple, efficient scheme that captures the essential behavior.
@@
float C = P.coat_weight; if (C <= 0.f) return;
@@
vec3 base_color = P.base_weight * P.base_color;
@@
vec3 Sc = P.subsurface_color;
@@
vec3 E_dielec = mix(mix(base_color, Sc, S), vec3(1.f-F0), T); // dielectric base albedo
vec3 E_metal = clamp(base_color * xi, 0.f, 1.f); // metallic base albedo
@@
vec3 Delta = max(1.f - K, 0.f) / (WHITE - E_base*K*$C$); // darkening factor, Eqn.76
@@
// Apply albedo scaling to model the layering of the coat on the base.
// The coat BSDF has the coat_weight presence weight functioning as a multiplier:
@@
// Account for absorption along the refracted entry/exit paths in the coat medium:
float mu_o = dot(sg.wo, P.coat_normal); mu_o = sqrt(1 - (1 - sqr(mu_o))/sqr(n_coat));
float mu_i = dot(sg.wi, P.coat_normal); mu_i = sqrt(1 - (1 - sqr(mu_i))/sqr(n_coat));
vec3 absorb_tint = pow($C$, 0.5f*(1.f/mu_i + 1.f/mu_o));
// The base BSDFs are then multiplied by the coat_throughput factor,
// which includes the base_darkening effect and the absorption tinting:
@@
vec3 coat_throughput = mix(WHITE, base_darkening * absorb_tint * (WHITE - coat_albedo), C);